Given logb2=5 and logb3=7, find logb144
Answer by lwsshak3(6472)

Given logb2=5 and logb3=7, find logb144
logb(144)
=logb(9*16)
=logb(3^2)+logb(2^4)
=2logb(3)+4logb(2)
=2*7+4*5
=14+20
logb(144)=34